OpenStudy (anonymous):

PLEASE HELP ASAP Will medal for correct answer! Eustis purchased a living room set for $2,784 using a 12-month deferred payment plan. The interest rate after the introductory period is 24.90%. A down payment of $225 is required as well as a minimum monthly payment of $107. What is the balance after the introductory period if only the minimum payment is made until then? A. $1,990.18 B. $2,053.07 C. $3,167.18 D. $3,274.18

OpenStudy (anonymous):

a

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@VeritasVosLiberabit

OpenStudy (anonymous):

b

OpenStudy (anonymous):

elaborate please

OpenStudy (anonymous):

got to love the people @ikatouni @jackieflawless who post answers with no work behind it

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@VeritasVosLiberabit

OpenStudy (anonymous):

@gdog361 trying to find out how to calculate deferment. Still looking. I was able to use a calculator and see that @ikatouni is right

OpenStudy (anonymous):

Let me try to find the formula

OpenStudy (anonymous):

With every formula I keep getting 745 \[P=\frac{ r(PV) }{ 1-(1+r)^{-n} }\]

OpenStudy (anonymous):

The interested accrued during deferment should be $778 which is added to the starting balance $2784

OpenStudy (anonymous):

then multiply 107*12+225 = 1509 and subtract this value from to obtain the balance after the introductory period $2053

OpenStudy (anonymous):

There is a step missing with the accrued value but hopefully this will be somewhat helpful. I willkeep looking and let you know if I find how the computer calculates $778 as the accrued interest

OpenStudy (anonymous):

\[P=\frac{ .249(2784) }{ 1-(1+.249)^{-12} }\]=745 2784+745=3529 107*12+225=1509 3529-1509=$2020

OpenStudy (anonymous):

r is rate, PV is stating value, n is the number of periods

OpenStudy (anonymous):

i figured it out and @jackieflawless was right, its A, they only want the balance after the introductory period

OpenStudy (anonymous):

its B

OpenStudy (anonymous):

i put a and got credit